OpenStudy (anonymous):

Sum and difference identities : prove cos3x=cos^3x-3sin^2xcosx

OpenStudy (anonymous):

I understand how to prove most of these but the exponents are throwing me off , could someone walk me through step by step please? Thank you!

OpenStudy (welshfella):

try usong cos (2 + x) = cos2x cos x - sin2x sinx

OpenStudy (welshfella):

= (2cos^2 x - 1) cos x - 2sinx cos x sin x

OpenStudy (welshfella):

= 2 cos^3 x - cosx - 2 sin^2x cosx

OpenStudy (welshfella):

hmm - doesn''t look like we are going anywhere with this

OpenStudy (freckles):

@welshfella it looks good to me now you need double angle identities

OpenStudy (freckles):

oh you did that I didn't read all the posts

OpenStudy (freckles):

\[(\cos^2(x)-\sin^2(x))\cos(x)-2\sin(x)\cos(x)\sin(x)\]

OpenStudy (freckles):

@bastiennecroix can you try to take it from here

OpenStudy (mathmale):

I originally shied away from this problem because of the ambiguity in your problem statement: cos3x=cos^3x-3sin^2xcosx. Does your "cos 3x" mean just that, or does it mean (cos x)^3? It may be worth your while to learn how to use Equation Editor, or at least the Draw utility, below: \[\cos 3x=\cos^3x-3\sin^2xcos x\]

OpenStudy (mathmale):

Please verify that this re-typed statement is correct (or not).

OpenStudy (michele_laino):

hint: if we apply the formula of Euler, we can write: \[\Large {e^{ix}} = \cos x + i\sin x\] from which we get: \[\Large {e^{i3x}} = \cos \left( {3x} \right) + i\sin \left( {3x} \right)\] furthermore, we have: \[\Large {e^{i3x}} = {\left( {{e^{ix}}} \right)^3} = {\left( {\cos x + i\sin x} \right)^3}\] so we get thie equality: \[\Large \cos \left( {3x} \right) + i\sin \left( {3x} \right) = {\left( {\cos x + i\sin x} \right)^3}\] please develop the third power at the right side, and equate the real parts of both sides

OpenStudy (welshfella):

ah - yes that might be the way to go = cos^3 x - sin^2 x cosx - 2 sin^2 x cos x

OpenStudy (welshfella):

thats it I picked the wrong identity for cos 2x

OpenStudy (mathmale):

You may want to use the sum formula for the cosine: cos (x+y) = cos x cos y - sin x sin y

OpenStudy (mathmale):

...so that cos 3x = cos (x + 2x)= ??? Please expand this. Decide for yourself whether your result will help you answer your original question.